# Bone: Block Affine Adaptation of Large Language Models

### v4

@@ Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. This paper introduces a novel PEFT technique distinct from LoRA, called Block-Affine Adaptation (Bone). By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants. Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation. Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants. Inspired by Pissa, we further proposed the “Weight Guide” theory to better utilize the information from the original weights. By integrating “Weight Guide” with Bone, we developed a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
@@ Results
LoRA 0.2365% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
Pissa 0.2365% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
Bone 0.2360% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
@@ 4.4 Bone vs Bat
Table 4: Comparing Bat-row, Bat-col on math tasks
@@ 5 Resource and efficiency
Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.95}$ kt/s

### v5

# Bone: Block-Affine Adaptation of Large Language Models

@@ Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. Many LoRA variants have emerged, yet they lack a design tailored to the characteristics of LLM weights and fail to leverage the original weights effectively. To address the sparsity of LLM weights, and drawing inspiration from GQA and MQA, we propose Block-Affine Adaptation (Bone), a novel PEFT technique distinct from LoRA. By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants. Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation. Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants. Inspired by Pissa, we propose a new theory called “Weight Guide” to better utilize the information embedded in the original weights. This approach extracts valuable information through a linear transformation of the original weight matrix using a trainable matrix. To validate the effectiveness of “Weight Guide” we combined it with Bone to create a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
@@ Results
LoRA 0.236% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
Pissa 0.236% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
Bone 0.236% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
@@ 4.4 Bone vs Bat
Table 4: Comparing Bone, Bat on math tasks
@@ 5 Resource and efficiency
Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.99}$ kt/s